Show solution.

How long can a car with a 12V battery of 60 A h capacity have the lights, of total resistance of 4 ohms, on before the battery runs down?


1
Expert's answer
2021-03-23T11:09:57-0400

Convert the voltage and capacity of the battery into energy units: 12V, 60A·h are


"E=12·60·3600=2592·10^3\text{ J}."

The power the lights consume is


"P=\frac{V^2}{R}."

The energy, on the other hand, is


"E=Pt,\\\space\\
t=\frac EP=\frac{ER}{V^2}=72000\text{ s},"

or 20 h.